# Blockchain Amplification Attack

### v2

@@ Abstract
Strategies related to the blockchain concept of Extractable Value (MEV/BEV), such as arbitrage, front- or back-running create an economic incentive for network nodes to reduce latency. A modified node, that minimizes transaction validation time and neglects to filter invalid transactions in the Ethereum P2P network, introduces a novel attack vector—Blockchain Amplification Attack. An attacker exploits those modified nodes to amplify an invalid transaction thousands of times, posing a threat to the entire network. To illustrate attack feasibility and practicality in the current mainnet, we 1) identify thousands of similar attacks in the wild, 2) mathematically model propagation mechanism, 3) empirically measure model parameters from our two monitoring nodes, and 4) compare performance with existing Denial-of-Service attacks through local simulation. We show that an attacker can amplify network traffic at modified nodes by a factor of 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack. Despite these risks, aggressive latency reduction may still be profitable enough to justify the existence of modified nodes. To assess this trade-off, we 1) simulate the transaction validation process in the local network and 2) empirically measure the latency reduction by deploying our modified node in the Ethereum testnet. We conclude with a cost-benefit analysis of skipping validation and provide mitigation strategies against this attack.
1 Introduction
The journey of transactions in blockchain P2P networks en route to their ultimate validation remains understudied. The specifics of handling pending transactions vary according to the client’s software version and its unique implementation. Under the concept of Extractable Value (MEV/BEV) [9, 45]—obtaining extra profits by reordering transactions, centralized business entities have the incentive to reduce latency to deliver pending transactions. Being the fastest in this space (as little as 1 millisecond) attracts users, bots, and validators who profit from arbitrage or front/back running. This scenario bears a resemblance to the traditional stock market where trading firms physically shortened the network paths to stock exchange matching engines and optimized networking configuration for quicker trade execution [34]. Recently, we notice certain modified nodes tailor their configurations, sidestep transaction validation processes to shorten latency, and introduce invalid transactions that are not supposed to persist in the network.
@@ 1 Introduction
Figure 1: Overview of the lockchain Amplification Attack, illustrating how one invalid transaction amplifies in the network.
@@
We formalize the Blockchain Amplification Attack. In Figure 1, the adversary sends an invalid transaction (a red arrow) to the modified node(s) that collectively propagate an invalid transaction (blue arrows) to the rest of the network. In this setting, the attacker not only disables the modified nodes and causes economic damages through excessive traffic, but also spreads invalid transactions to slow down the entire network. This paper fully describes such potential attack opportunities, documents the prevalence of similar attacks in the wild, empirically/simulatively models the network damage, and quantifies the trade-offs associated with skipping validations. Due to the absence of penalty mechanism in the current P2P network, we demonstrate that the attack is practical, cost-effective, and a threat to the whole network.
@@
We first mathematically model the Blockchain Amplification Attack and pinpoint the parameters that affect how an attacker’s invalid transactions can get amplified in the whole network (i.e., amplification factor). While considering the specifics of each software/version on transaction forwarding, our model establishes a framework to quantify the increase in both the amount and the economic cost of egress traffic at the modified nodes.
@@
We next demonstrate the attack feasibility in practice. In particular, we analyze the set of pending transactions (called “txpool”) for centralized services. Some of them propagate transactions much faster than the others but deliver invalid transactions, suggesting the lack of validation checks. We identify 2,591 similar attack instances across 345 Ethereum blockchain addresses that target these services. We classify those instances based on the attacker’s observed strategies and estimate the attack cost, size, and intensity.
@@
To correctly estimate amplification factors (i.e., the attack impact) in the current mainnet, we design two customized monitoring Geth nodes (the most popular Ethereum node client) to infer the network topology. Our custom nodes scan the activities on the P2P layer and record specific types of messages before they undergo processing, resulting 2.5 billion observations over 5 months. This infrastructure allows us to propose a new method for inferring the number of active peer connections, which serves as a cost-effective/ethical alternative to previously proposed methods [13, 10, 4, 35]. We further show that 1.5% of nodes in our dataset exhibit lenient transaction validation (“modified nodes”). Many of those nodes appear to come from the same entity given the identical git commit in a node name, which we do not find in any public software repository. We show that the attacker can amplify outgoing network traffic at modified nodes by a factor of at least 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack.
@@
In addition, we conduct attack simulations on the local P2P network and confirm that our proposed attack can evict as many honest transactions (from both the txpool/the block) as the existing DoS attacks but at significantly lower costs.
@@
Furthermore, we experiment to quantify the benefits of skipping transaction validation, specifically the amount of time saved for each validation process. We first fork our archive node (the node stores all the historical states of the blockchain) to simulate/dissect the process of Geth’s transaction validation in the locally controlled environment. We find that validating one transaction takes roughly 1 millisecond and checking account status (nonce/balance) is the most time-consuming (86%). Moreover, we empirically measure the latency reduction by deploying the modified/regular nodes in the Ethereum testnet for 14 weeks.
Finally, we perform a benefit-cost analysis of a modified node based on time-to-money conversion [52, 46] and offer three possible mitigation tailored to Ethereum: 1) enforcing a stricter txpool policy, 2) postponing the validation process, and 3) introducing a node reputation system.
@@
  • We mathematically define the Blockchain Amplification Attack and show the attack feasibility based on similar attacks in the wild.
  • By designing and deploying our custom monitoring nodes, we infer the Ethereum P2P network topology and illustrate the attack efficiency.
  • Simulations on the local network confirm the superiority of our attack over the existing DoS attacks.
  • We simulate, dissect and empirically measure the transaction validation process for latency reduction.
2 Background
@@
2.1 Blockchain and P2P network
Three aspects of the blockchain ecosystem are particularly pertinent to our paper: 1) the formation of a P2P network by nodes, 2) the validations of transactions, and 3) the interaction with users outside the P2P network.
Ethereum nodes use the Kademlia [39] distributed hash table for connectivity. When a node launches, it initially connects to hardcoded bootstrap nodes, which provide a list of potential peers. The node identifier, known as “enode,” comprises the node ID (randomly generated from the node’s public key), IP address, and port number. The node calculates the distance between possible peers based on node IDs, divides them into buckets, and populates the list of peers, [^1] The node calculates the XOR of two node IDs: Keccak256(node1’s id) XOR Keccak256(node2’s id) as a distance, and group the possible peers into 256 buckets: for each $i$ (bucket), where $2^{i}\leq distance<2^{i+1}$) making peer selection arbitrary [16]. Geth by default allows 50 connections, while Erigon and Nethermind permit 100. Geth allocates one-third of connections to active peer searching and uses the other two-thirds for passively accepting inbound connections. Ethereum’s mainnet shares the same underlying P2P network with other chains (e.g., Ethereum Classic or Testnets such as Holesky). Geth nodes immediately disconnect from peers on different chains.
After establishing a connection, nodes initiate the exchange of messages to synchronize blockchain information (blocks and transactions), following the Ethereum Wire Protocol [15]. For every transaction, a node must validate the integrity of each transaction, which contains several parameters, regardless of encoding type. Ethereum is an account-based ledger, meaning each address maintains the balance and the transaction index called a “nonce” (starting from 0, incrementing per transaction). The amount of “gas” represents the computational costs of the transaction. We multiply by the gas price to obtain the total transaction fees. The transaction is either to 1) send “value” to another address, or 2) execute a smart contract where “data field” is used to specify the function and its input. The node checks the following conditions for each transaction:
@@ 2.1 Blockchain and P2P network
Beyond protocol-level verifications, each software incorporates supplementary checks to mitigate DoS (Denial-of-Service) attacks. For example, Geth employs the “gas bump rule,” requiring the gas fee for a transaction with the same sender and nonce to be higher than the original transaction with a specified increase (e.g., 10% in Geth). The node should not accept replicated transactions with the same address, nonce, and gas price. Li et al. [35, §5.1] or Yaish et al.  [55, §4.1] provide detailed explanations.
If the transaction is valid, it gets added to the buffer referred to as “txpool” (or mempool, tx-queue); otherwise, it should be dropped. Each software implementation imposes its own limit on the number of transactions accepted in the txpool; legitimate transactions may be discarded if the txpool reaches its maximum capacity.
When a new transaction enters the node’s txpool, the node broadcasts it to the rest of the network using two message options: 1) broadcast Transactions (0x02) – RLP (Recursive-Length Prefix)-encoded raw transactions that include all parameters, or 2) announcement NewPooledTransactionHashes (0x08) – transmitting only transaction hashes. Typically, the node broadcasts Transactions (0x02) to a small subset of nodes for efficiency, and announces NewPooledTransactionHashes (0x08) to the remaining nodes. If the node lacks information about a transaction based on its hash (0x08), it requests a peer to send transaction content via GetPooledTransactions (0x09), and the recipient node responds with PooledTransactions (0x0a). We describe software-specific implementations in §4.3.
@@
When users fetch the blockchain data or initiate a transaction, they typically interact with a P2P node through Remote Procedure Call (RPC) requests, and nodes execute these requests. Various publicly accessible RPC endpoints (e.g., Infura) allow users (rate limited) free access (including pending transactions in the txpool). Txpool view providers (e.g., bloXroute, Eden, Chainbound), in particular, offer a “better” view of txpool in terms of latency. For example, bloXroute employs a Blockchain Distribution Network (BDN) infrastructure to improve blockchain scalability [32, 3]. The globally distributed nodes index transactions to reduce the size, internally propagate without validations (“cut-through routing”), and employ dynamic routing to overcome network overhead.
2.2 Security in P2P network
@@
While a P2P network serves as the foundation of decentralized blockchains, sharing the same information among all peers is challenging, and often results in network forks [12]. Forks expose the network to double-spending [30, 22], eclipse [27] or selfish mining [19] attacks that all partition the network to block information flow. In the PoW blockchain, Luu et al. [37] proposed the “Verifier’s dilemma,” wherein rational miners opt to forgo block validation to allocate more time to PoW mining. Das et al.[11] demonstrated that PoW miners who bypass validation stand a significantly higher chance of winning the block compared to their mining power. The same applies to transaction latency especially if the blockchain supports decentralized exchanges, which allow transaction re-ordering attacks [9, 45]. As long as there is an economic incentive to capture transactions fast, centralized entities strive to minimize latency, which might make the P2P network more centralized and introduce new security vulnerabilities.
@@
Our proposed attack stems from not only DoS attack [33] but also Economic Denial of Sustainability (EDoS) attack. Hoff [29] first presented the EDoS concept in a 2008 blog post. EDoS, akin to a traditional Denial of Service (DoS) attack, focuses on inflicting financial losses by targeting the victim’s traffic usage [53]. Attackers exploit the knowledge of a victim server hosted on a cloud service by manipulating traffic usage to inflate bills, making the victim’s operation economically unsustainable. The same issue applies to blockchain P2P networks where many nodes have public IP addresses, and use cloud platforms such as AWS.
3 Data
@@
We use publicly available data (RPC nodes, txpool providers) from Flashbots [25], as well as measurement through our nodes deployed in a P2P network (Figure 2).
3.1 Public txpool data
@@
Each node has a different view of the txpool, meaning that they hold a different set of pending transactions. A node that opens up its API connection to the public makes its txpool accessible (newPendingTransactions). Flashbots [^3] https://www.flashbots.net/ has been providing publicly available txpool information on “Mempool Dumpster” [25] every day, and we use its data from September 1st, 2023 to January 11th, 2024. Flashbots collects a set of pending transactions from 1) RPC providers and generic nodes – Infura, Alchemy, A-pool, [^4] A regular network node with the optimized peering setting whom Flashbots listens to. Flashbots’ local node, Mempoolguru [1], and 2) the infrastructure txpool providers – bloXroute, Chainbound, Eden. All the transactions contain transaction hash, source (i.e., which service/node), and observation timestamp (in milliseconds). Some portions of transactions include the details of transactions (e.g., sender, receiver, gas, value, and data fields). We remove repetitions (identical transactions) from the same source to avoid double-counting. Consequently, our summary statistics slightly diverge from Flashbots’.
3.2 Private monitoring node data
Figure 2: An overview of our monitoring node.

### v3

@@ Abstract.
Strategies related to the blockchain concept of Extractable Value (MEV/BEV), such as arbitrage, front-, or back-running create strong economic incentives for network nodes to reduce latency. Modified nodes, that minimize transaction validation time and neglect to filter invalid transactions in the Ethereum peer-to-peer (P2P) network, introduce a novel attack vector—a Blockchain Amplification Attack. An attacker can exploit those modified nodes to amplify invalid transactions thousands of times, posing a security threat to the entire network. To illustrate attack feasibility and practicality in the current Ethereum network (“mainnet”), we 1) identify thousands of similar attacks in the wild, 2) mathematically model the propagation mechanism, 3) empirically measure model parameters from our monitoring nodes, and 4) compare the performance with other existing Denial-of-Service attacks through local simulation. We show that an attacker can amplify network traffic at modified nodes by a factor of 3,600, and cause economic damages of approximately 13,800 times the amount needed to carry out the attack. Despite these risks, aggressive latency reduction may still be profitable enough for various providers to justify the existence of modified nodes. To assess this trade-off, we 1) simulate the transaction validation process in a local network and 2) empirically measure the latency reduction by deploying our modified node in the Ethereum test network (“testnet”). We conclude with a cost-benefit analysis of skipping validation and provide mitigation strategies against the blockchain amplification attack.
Blockchain, Security, DoS, P2P network, Ethereum
copyright: ccjournal: POMACSjournalyear: 2025journalvolume: 9journalnumber: 1article: 4publicationmonth: 3doi: 10.1145/3711697ccs: Security and privacy Denial-of-service attacksccs: Networks Network measurement
1. Introduction
In “Flash Boys” (Lewis, 2014), Lewis famously describes how, in the traditional stock market, lower network latency to stock exchange matching engines provides such a competitive advantage that trading firms have started to physically shorten network links – using, e.g., dedicated fiber – and routinely optimize network configuration to reduce latency.
Could the same type of optimizations apply to decentralized finance (DeFi), i.e., in the context of blockchain peer-to-peer (P2P) networks? In principle, yes: the concept of Extractable Value (MEV/BEV) (Daian et al., 2020; Qin et al., 2022), i.e., maximizing profit by reordering transactions, implies that centralized business entities have an incentive to reduce latency to deliver pending transactions. Being the fastest to get transactions included in a block – even if just by milliseconds – can attract users, bots, and validators who can profit from arbitrage, front- or back-running, and sandwich attacks (Zhou et al., 2021b).
@@ 1. Introduction
Figure 1. Overview of the Blockchain Amplification Attack, illustrating how one invalid transaction amplifies in the network.
@@
As a case in point, we recently notice that certain modified nodes in the Ethereum network tailor their configurations, sidestep transaction validation processes to shorten latency, which introduce invalid transactions that are not supposed to propagate in the network. Unfortunately, these optimizations open the door to a Blockchain Amplification Attack, which we formalize and study in this paper. In this attack, depicted in Figure 1, 1) the adversary $A$ sends an invalid transaction (denoted by the red arrow) to a modified node $M$, that 2) shares it with other modified nodes (black arrows), which will 3) collectively further propagate it (blue arrows) to the rest of the network, amplifying the effect of the invalid transaction on the network. Indeed, the attacker can not only hamper modified nodes (and cause economic damage through excessive traffic), but, by also spread invalid transactions to slow down the entire network. This paper fully describes such attack opportunities, documents the prevalence of similar attacks in the wild, models network damage empirically and by simulation, and quantifies the trade-offs associated with skipping validations. Due to the absence of a penalty mechanism for sending invalid transactions, we demonstrate that the attack is feasible, impactful, and cost-effective in the current Ethereum P2P network (“mainnet”).
@@
We first mathematically model the transaction propagation mechanism and pinpoint the parameters that affect how an attacker’s invalid transactions can get amplified in the whole network (i.e., the amplification factor). While considering the specifics of each software/version on transaction forwarding, our model establishes a framework to quantify the increase in both the amount and the economic cost of invalid transactions.
@@
We next demonstrate the attack feasibility in the wild. In particular, we analyze the set of pending transactions (usually referred to as “txpool”) for centralized entities that deliver transactions to users. Some propagate transactions much faster than the others but propagate invalid transactions, suggesting the lack of validation checks. We identify 2,591 similar attack instances across 345 Ethereum addresses that target these entities. We classify those instances based on the attacker’s observed strategies and estimate the attack cost, size, and intensity.
@@
To correctly estimate amplification factors (i.e., the impact of the attack) in the current mainnet, we design two customized monitoring Geth nodes (the most popular Ethereum node client) to infer the network topology. Our custom nodes scan the activities on the P2P layer and record specific types of messages before they undergo processing, resulting 2.5 billion observations. Our infrastructure allows us to perform a new method for inferring the number of active peer connections—a cost-effective and ethical alternative to previously proposed methods (Delgado-Segura et al., 2019; Daniel et al., 2019; Cao et al., 2020; Li et al., 2021a). We further show that 1.5% of nodes exhibit lenient transaction validation (“modified nodes”). Looking at the git commit hash of the modified nodes, which shows unique and non-public properties of the software, many of these nodes appear to originate from the same entity. All in all, we show that the attacker can amplify outgoing network traffic at modified nodes by a factor of at least 3,600, and cause economic damages 13,800 times the amount needed to carry out the attack.
@@
In addition, we conduct attack simulations in a local network and confirm that our proposed attack can evict as many honest transactions (from both the txpool and the block) as other proposed DoS attacks but at significantly lower costs.
@@
Furthermore, we experiment to quantify the benefits of skipping transaction validation, specifically the amount of time saved for each validation process. We first simulate the process of Geth’s transaction validation in a locally controlled environment. We find that validating one transaction takes roughly 1 millisecond and checking account status (nonce/balance) is the most time-consuming (86% of total validation time). We also empirically measure the latency reduction by deploying the modified/regular nodes in the Ethereum test network (“testnet”) for 14 weeks.
Based on the latency gains from skipping validations, we estimate the maximum economic gains (0.00002 ETH $\approx$ \$0.050 per millisecond (Wahrstätter et al., 2023; Schwarz-Schilling et al., 2023; Babel et al., 2024)), and compare it with the potential attack damage. We finally offer three possible mitigations: 1) enforcing a stricter txpool policy, 2) postponing the validation process, and 3) introducing a node reputation system.
@@
  • We mathematically define a Blockchain Amplification Attack and show attack feasibility based on similar attacks found in the wild.
  • By designing and deploying our custom monitoring nodes, we infer the Ethereum P2P network topology and illustrate the efficiency of the attack.
  • Simulations on the local network confirm that the blockchain amplification attack is more cost-effective for the attacker than other existing DoS attacks.
  • We simulate and empirically measure the transaction validation process for latency reduction.
2. Background
@@
2.1. Blockchain and P2P network
Three aspects of the blockchain ecosystem are particularly pertinent to our paper: 1) P2P network formation, 2) transaction validation, and 3) interactions with users outside the P2P network.
Ethereum nodes use the Kademlia (Maymounkov and Mazieres, 2002) distributed hash table for connectivity. When a node launches, it initially connects to hardcoded bootstrap nodes, which provide a list of potential peers. The node identifier, known as “enode,” comprises the node ID (randomly generated from the node’s public key), IP address, and port number. The node calculates the distance between possible peers based on node IDs, divides them into buckets, and populates the list of peers, [^1] The node calculates the XOR of two node IDs: Keccak256(node1’s id) XOR Keccak256(node2’s id) as a distance, and group the possible peers into 256 buckets: for each $i$ (bucket), where $2^{i}\leq distance<2^{i+1}$) making peer selection arbitrary (Ethereum, 2023b). Geth by default allows 50 connections, while Erigon and Nethermind permit 100. Geth allocates one-third of connections to active peer searching and uses the other two-thirds for passively accepting inbound connections. Ethereum’s mainnet shares the same underlying P2P network with other chains (e.g., Ethereum Classic or Testnets such as Holesky). Geth nodes immediately disconnect from peers on different chains.
After establishing a connection, nodes initiate the exchange of messages to synchronize blockchain information (blocks and transactions), following the Ethereum Wire Protocol (Ethereum, 2023a). For every transaction, a node must validate the integrity of each transaction, which contains several parameters, regardless of encoding type. Ethereum is an account-based ledger, meaning each address maintains the balance and the transaction index called a “nonce” (starting from 0, incrementing per transaction). The amount of “gas” represents the computational costs of the transaction. We multiply by the gas price to obtain the total transaction fees. The transaction is either to 1) send “value” to another address, or 2) execute a smart contract where “data field” is used to specify the function and its input. The node checks the following conditions for each transaction:
@@ 2.1. Blockchain and P2P network
Beyond protocol-level verifications, each software incorporates supplementary checks to mitigate DoS (Denial-of-Service) attacks. For example, Geth employs the “gas bump rule,” requiring the gas fee for a transaction with the same sender and nonce to be higher than the original transaction with a specified increase (e.g., 10% in Geth). The node should not accept replicated transactions with the same address, nonce, and gas price. Li et al. (Li et al., 2021a, §5.1) or Yaish et al.  (Yaish et al., 2023, §4.1) provide detailed explanations.
If the transaction is valid, it gets added to the buffer referred to as “txpool” (or mempool, tx-queue); otherwise, it should be dropped. Each software implementation imposes its own limit on the number of transactions accepted in the txpool; even legitimate transactions may be discarded if the txpool reaches its maximum capacity.
When a new transaction enters the node’s txpool, the node broadcasts it to the rest of the network using two message options: 1) broadcast Transactions (0x02) – RLP (Recursive-Length Prefix)-encoded raw transactions that include all parameters, or 2) announce NewPooledTransactionHashes (0x08) – transmitting only transaction hashes. Typically, the node broadcasts 0x02 to a small subset of nodes for efficiency, and announces 0x08 to the remaining nodes. If the node lacks information about a transaction based on its hash (0x08), it requests a peer to send transaction content via GetPooledTransactions (0x09), and the recipient node responds with PooledTransactions (0x0a). We describe software-specific implementations in §4.3.
@@
When users fetch the blockchain data or initiate a transaction, they typically interact with a P2P node through Remote Procedure Call (RPC) requests, and nodes execute these requests. Given the high cost of maintaining own node, there exist centralized entities (e.g., Infura, Alchemy) that open up their RPC endpoints or direct node connections for users. Some (e.g., bloXroute, Eden, Chainbound) optimize the network configurations and deliver pending transactions faster than other nodes as a service. For example, bloXroute deploys nodes across globe, internally indexes transactions, propagates without block validations (“cut-through routing”), and employs dynamic routing (Klarman et al., 2018; Bloxroute, 2019). Users sometimes pay subscription fees to connect to those special nodes.
2.2. Security in P2P network
@@
While a P2P network serves as the foundation of decentralized blockchains, sharing the same information among all peers is challenging, and often results in network forks (Decker and Wattenhofer, 2013). Forks expose the network to double-spending (Karame et al., 2012; Gervais et al., 2016), eclipse (Heilman et al., 2015) or selfish mining (Eyal and Sirer, 2018) attacks that all partition the network to block information flow. In the PoW (Proof of Work) blockchain, Luu et al. (Luu et al., 2015) proposed the “Verifier’s dilemma,” wherein rational miners opt to forgo block validations to allocate more time to PoW mining. Das et al.(Das et al., 2021) demonstrated that PoW miners who bypass validation stand a significantly higher chance of winning the block compared to their mining power. Despite the similarity in concept, our work focuses on transaction validations in the P2P network. In our scenario, the centralized entities (i.e., validators, relays) forward transactions without validations, giving users extra time to execute transaction re-ordering attacks (Daian et al., 2020; Qin et al., 2022). As long as there is an economic incentive to capture transactions fast, centralized entities strive to minimize latency, which might make the P2P network more centralized and introduce new security vulnerabilities.
@@
Our proposed attack stems from not only DoS attack (Kumar, 2007) but also Economic Denial of Sustainability (EDoS) attack. Hoff (Hoff, 2008) first presented the EDoS concept in a 2008 blog post. EDoS, akin to a traditional Denial of Service (DoS) attack, focuses on inflicting financial losses by targeting the victim’s traffic usage (Wang et al., 2016). Attackers exploit the knowledge of a victim server hosted on a cloud service by manipulating traffic usage to inflate bills, making the victim’s operation economically unsustainable. The same issue applies to blockchain P2P networks where many nodes have public IP addresses, and use cloud platforms such as AWS.
3. Data
@@
We use publicly available data (RPC nodes, txpool providers) from Flashbots (Hager, 2023), as well as measurement through our nodes deployed in a P2P network (Figure 2).
3.1. Public txpool data
@@
Each node has a different view of the txpool, meaning that they hold a different set of pending transactions. A node that opens up its API connection to the public makes its txpool accessible (newPendingTransactions). Flashbots [^3] https://www.flashbots.net/ has been providing publicly available txpool information on “Mempool Dumpster” (Hager, 2023) every day, and we use its data from September 1st, 2023 to January 11th, 2024. Flashbots collects a set of pending transactions from 1) RPC providers and generic nodes – Infura, Alchemy, A-pool, [^4] A regular network node with the optimized peering setting whom Flashbots listens to. Flashbots’ local node, Mempoolguru (mem, 2023), and 2) the infrastructure txpool providers – bloXroute, Chainbound, Eden. All the transactions contain transaction hash, source (i.e., which service/node), and observation timestamp (in milliseconds). Some portions of transactions include the details of transactions (e.g., sender, receiver, gas, value, and data fields). We remove repetitions (identical transactions) over time to avoid double-counting. Consequently, our summary statistics slightly diverge from Flashbots’.
3.2. Private monitoring node data